OpenStudy (anonymous):

Which number is divisible by 3 and 5? A. 130 B.325 C.4,113 D.2,040

OpenStudy (zzr0ck3r):

if its divisible by 3, then the sum of its digits is divisible by 3 if its divisible by 5, then the last digit is 0 or 5

OpenStudy (anonymous):

I don't understand. Please explain?

OpenStudy (zzr0ck3r):

take the first one 130 it is divisible by 5, because the last digit is 0, but if you add 1+3=4 is not divisible by 3 so the number is also not for 325 again its divisible by 5, but 3+2+5 = 10 is not divisible by 3 so neither is the number for 4113 this is not divisible by 5 because the last digit is NOT 0 or 5, its 3 for 2040 this is divisible by 5, and 2+0+4+0 = 6 is divisible by 3, so 2040 is divisible by 3 and 5

OpenStudy (zzr0ck3r):

is 3123420 divisible by 5 and 3?

OpenStudy (anonymous):

no.